OpenStudy (anonymous):

What is the 8th term in the pattern 2, 10, 18, 26, ...? 52 54 56 58

OpenStudy (anonymous):

58

OpenStudy (the_fizicx99):

It's increasing by 8, So you can just use a explicit formula: f(n) = f(1) + d(n - 1) Find the 8th term, f(8) = f(1) + 8(8-1) f(8) = f(1) + 8(7) f(8) = f(1) + 56 f(8) = 2 + 56 f(8) = 58

OpenStudy (the_fizicx99):

@collegekids please refrain from giving out answers

OpenStudy (anonymous):

lol i was just adding 8 on the calculator

OpenStudy (the_fizicx99):

.___.

OpenStudy (anonymous):

how hard is that

OpenStudy (the_fizicx99):

>.< she needs to learn the point and how to do it.

OpenStudy (anonymous):

im bad at explaining lol, but great at math